OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that has a slope of -2 and passes through the point (2, -8). y = -2x - 4 y = -2x - 8 y = -2x + 18 y = -2x – 12

OpenStudy (anonymous):

y-y1=m(x-x1) y-(-8)=-2(x-2) y+8=-2x+4 y=-2x-4

OpenStudy (anonymous):

A dog owner buys a 50 lbs bag of dog food. Each day she used 2 lbs of dog food to feed her dogs. After 15 days how much dog food remains? 20 24 35 48

OpenStudy (anonymous):

20